Chemistry. E-Chem Testing has a solution that is 80% base and another that is 30% base. A technician needs 150 L of a solution that is 62% base. The 150 L will be prepared by mixing the two solutions on hand. How much of each should be used?
Answer by mananth(16946)

percent ---------------- quantity
Base I 80 ---------------- x L
Base II 30 ------ 150 - x L
Mixture 62.00% ---------------- 150

80x+30(150-x)=62*150

80x+4500-30x=9300
80x-30x =9300-4500
50x=4800
/ 50
x =96L 80.00% Base I
54L 30.00% Base II

m.ananth@hotmail.ca